solve 2x^2 +13x+ 20 = 0 using complete the square
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

solve 2x^2 +13x+ 20 = 0 using complete the square
-----------
2x^2 +13x+ 20 = 0
x^2 + 6.5x + 10 = 0
x^2 + 6.5x = -10
x^2 + 6.5x + (6.5/2)^2 = -10 + (6.5/2)^2 = 0.5625
(x + 3.25)^2 = 0.75^2
x + 3.25 = 0.75
x = -2.5
===============
x + 3.25 = -0.75
x = -4

Answer by ikleyn(53763)

.
solve 2x^2 +13x+ 20 = 0 using complete the square
~~~~~~~~~~~~~~~~~~~~


The correct solution is THIS:

2x^2 +13x+ 20 = 0

x^2 + 6.5x + 10 = 0

x^2 + 6.5x + (6.5/2)^2 = -10 + (6.5/2)^2 = 0.5625

(x + 3.25)^2 = 0.75^2

x + 3.25 = +/- 0.75

x = -3.25 +/- 0.75


There are TWO solutions:  x= -3.25 + 0.75 = -2.5,   and   x= -3.25 - 0.75 = -4.


Check.  2*(-2.5)^2 +13*(-2.5) + 20 = 0   ! Correct !

        2*(-4)^2 +13*(-4) + 20 = 0       ! Correct ! 

Solved.